OpenStudy (anonymous):

A tourist being chased by an angry bear is running in a straight line to- ward his car at a speed of 4.0 m/s. The car is a distance d away. The bear is 26 m behind the tourist and running at 6.0 m/s. The tourist reaches the car safely. What is the maximum possible value for d? I cannot wrap my head around this one :P Thanks

OpenStudy (anonymous):

if they both cover the same distance (which will then be the max safe distance) 6t -26m = d = 4t so 2t =26m t is 13 d= 4*13 = 52